Study whether the vectors v1(1, 1, 2), v2(2, 3, 0), v3(0, 1, 2) in R 3 are linearly independent.


1
Expert's answer
2022-01-25T17:08:57-0500

A sequence of vectors "\vec v_1, \vec v_2,..., \vec v_k"  from a vector space "V" is said to be linearly dependent, if there exist scalars "a_1, a_2, ..., a_k" not all zero, such that


"a_1\vec v_1+a_2\vec v_1+...+a_k\vec v_k=\vec 0"

where "\vec 0" denotes the zero vector.

Consider the set of vectors "\vec v_1=(1,1,2), \vec v_2=(2,3,0), \vec v_3=(0,1,2)" then the condition for linear dependence seeks a set of non-zero scalars, such that


"\begin{bmatrix}
 1 & 2 & 0 \\
 1 & 3 & 1 \\
 2 & 0 & 2 \\
\end{bmatrix}\begin{bmatrix}
 a_1 \\
 a_2 \\
a_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
0
\end{bmatrix}"

Augmented matrix


"\begin{bmatrix}
 1 & 2 & 0 & \ 0 \\
 1 & 3 & 1 & \ 0 \\
 2 & 0 & 2 & \ 0 \\
\end{bmatrix}"

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & 2 & 0 & \ 0 \\
 0 & 1 & 1 & \ 0 \\
 2 & 0 & 2 & \ 0 \\
\end{bmatrix}"

"R_3=R_3-2R_1"


"\begin{bmatrix}
 1 & 2 & 0 & \ 0 \\
 0 & 1 & 1 & \ 0 \\
 0 & -4 & 2 & \ 0 \\
\end{bmatrix}"

"R_1=R_1-2R_2"


"\begin{bmatrix}
 1 & 0 & -2 & \ 0 \\
 0 & 1 & 1 & \ 0 \\
 0 & -4 & 2 & \ 0 \\
\end{bmatrix}"

"R_3=R_3+4R_2"


"\begin{bmatrix}
 1 & 0 & -2 & \ 0 \\
 0 & 1 & 1 & \ 0 \\
 0 & 0 & 6 & \ 0 \\
\end{bmatrix}"

"R_3=R_3\/6"


"\begin{bmatrix}
 1 & 0 & -2 & \ 0 \\
 0 & 1 & 1 & \ 0 \\
 0 & 0 & 1 & \ 0 \\
\end{bmatrix}"

"R_1=R_1+2R_3"


"\begin{bmatrix}
 1 & 0 & 0 & \ 0 \\
 0 & 1 & 1 & \ 0 \\
 0 & 0 & 1 & \ 0 \\
\end{bmatrix}"

"R_2=R_2-R_3"


"\begin{bmatrix}
 1 & 0 & 0 & \ 0 \\
 0 & 1 & 0 & \ 0 \\
 0 & 0 & 1 & \ 0 \\
\end{bmatrix}"

Then "a_1=a_2=a_3=0."

Therefore the vectors "\vec v_1=(1,1,2), \vec v_2=(2,3,0), \vec v_3=(0,1,2)" in "R ^3" are linearly independent.